I really need help understanding how to work this word problem out.
If there is a goat tied to a rectangular barn on a 50 foot lead and the barn is 20 feet by 20
feet (floor), what is the maximum grazing area? If there are regions you can't find the area of,
provide as good an estimate as you can. Assume the goat is tied to a corner outside the barn,
cannot get in, and that the barn is not grazing area. (Remember, this will be based on parts of
circles, no other shapes...the goat's rope will only get shorter when he tries to go around the
barn...)
6. How much of the 50 foot circle can the goat reach without getting interrupted by the barn?
What is that area?

Answer by edjones(8007)

I'll answer the 2nd question 1st.
It can graze 3/4 of a circle with a 50' radius.
That area = pi*50^2= 2500*pi*(3/4)= 1875*pi
Of the area not covered it has two 1/4 circles= 1 semicircle with a radius of 30' that area = pi*30^2*.5= 450*pi.
1875*pi + 450*pi = 2325*pi sq. ft.
Ed